### v1

# Dynamic Moiré-like pattern in non-Hermitian Wannier-Stark ladder system

Abstract
We study the dynamical behavior of the non-Hermitian Wannier-Stark ladder system, which is a non-Hermitian Su-Schrieffer-Heeger chain with a position-dependent real potential. In the presence of a linear external field, we employ the non-Hermitian Floquet method and find that the energy levels are sensitive to the field. The system exhibits two distinct dynamic behaviors separated by an exceptional point: one in the $\mathcal{PT}$ symmetrical region associated with two real Wannier-Stark ladders, and another in the $\mathcal{PT}$ symmetry-breaking region associated with complex conjugate ladders. As the boundary between the two regions, two ladders coalesce into a single ladder. In the case of a non-linear field, these two distinct regions appear alternately along the chain, exhibiting dynamic Moiré-like patterns.
I Introduction
Wannier-Stark (WS) ladders are a theoretical model in solid-state physics that describe the quantization of electronic states in a crystal under the influence of a constant electric field Bloch1929 ; Wannier1959 ; Wannier1960 ; Glueck2002 ; Waschke1993 . It is closely related to Bloch oscillations (BOs) due to the equidistant spectrum. In the past decade, such a phenomenon has attracted much attention in cold-atoms physics and photonics due to applications in interferometric measurements and as a method for manipulating localized wave packets Breid2006 ; Breid2007 ; Dreisow2009 ; Kling2010 ; Ploetz2011 . It can be simulated using artificial quantum systems, such as superconducting circuits Song2024 . On the other hand, WS ladders are an essential theoretical tool for understanding and exploring particle transport phenomena in quantum systems and have potential applications in superconducting quantum computing and quantum simulation. In the previous study, encompassing both theoretical and experimental aspects, only a linear external field was considered because it is central to the formation of an equidistant spectrum.
The main question we address in this work is how a weak non-linear field influences the dynamics of a WS system. Intuitively, such a field results in only a slight change in the oscillating frequency along the chain. Nevertheless, as will be shown, a weak non-linear field leads to a fascinating phenomenon when a non-Hermitian double-well lattice system is considered. In this work, we consider a non-Hermitian Su-Schrieffer-Heeger (SSH) chain with a position-dependent real potential, which is schematically illustrated in Fig. 1(a). In the simple case of zero non-Hermitian hopping ($\beta=0$) and a linear field, the energy levels can be easily obtained as two sets of WS ladders with $E_{\pm}/\omega=2m\omega$$\pm\sqrt{\left(J/\omega\right)^{2}+1/4}$, where $m$ is an integer. Obviously, the energy difference can be zero or $\omega$, corresponding to two identical or a single WS ladder, respectively, as $\omega$ varies. Importantly, such a dramatic change can be induced by a shift in the frequency, $\Delta\omega\approx\omega^{2}/8$. This indicates that such a transition is very sensitive to the value of $\omega$, especially in the limit of small $\omega$. The corresponding distinct dynamical behaviors can emerge alternatively along the chain when a weak non-linear field is applied. Furthermore, when $\beta$ is nonzero, the non-Hermitian term can amplify this phenomenon.
One of the unique features of a non-Hermitian system is the violation of conservation law of the Dirac probability, based on which, the complex potential and non-Hermitian hopping strength are employed to describe open systems phenomenologically J2004 . Furthermore, unconventional propagation of light associated with the gain/loss has been demonstrated by engineering effective non-Hermitian Hamiltonians in optical systems Guo2009 ; Makris2008 ; Musslimani2008 ; Klaiman2008 ; Longhi2009 ; ElGanainy2007 ; Zheng2010 .
Exceptional points (EPs), as an exclusive feature of a non-Hermitian system, are degeneracies of non-Hermitian operators Berry2004 ; Heiss2012 ; Miri2019 ; Zhang2020 . The corresponding eigenstates coalesce into one state resulting to the incompleteness of Hilbert space. The peculiar features around EP have sparked tremendous attention to the classical and quantum photonic systems Doppler2016 ; Xu2016 ; Assawaworrarit2017 ; Wiersig2014 ; Wiersig2016 ; Hodaei2017 ; Chen2017 .Notably, a coalescing state has an exclusive feature. Both theoretical and experimental works not only give an insight into the dynamical property of the non-Hermitian Hamiltonian but also provide a platform to implement the novel optical phenomenon.
In this work, we study the dynamical behavior of the non-Hermitian WS ladder system with a weak non-linear field. Analytical analysis and numerical simulation based on the non-Hermitian Floquet method reveal that a system with a linear potential exhibits the following features: (i) The spectrum of the system consists of two sets of complex WS ladders; (ii) they can be two real WS ladders with identical level spacing, a single coalescing WS ladders referred to as an EP ladder, and complex conjugate piar of WS ladders; (iii) As the boundary between the complex and real regions, the EP ladder appears periodically as the slope of the linear field changes slightly. These allow for the appearance of dynamic Moiré-like patterns in the system with a weak non-linear field.
In general, Moiré patterns emerge from the superposition of two periodic structures that have either slightly different periods or different orientations. These patterns have been realized in materials Yankowitz2012 ; Ponomarenko2013 ; Dean2013 ; Hunt2013 ; Woods2014 ; Nakajima2016 ; Lohse2015 . Recently, the influence of Moiré patterns in physical systems has attracted considerable interest Ponomarenko2013 ; Dean2013 ; Hunt2013 ; Gorbachev2014 ; Song2015 ; Jung2014 , particularly in non-Hermitian systems and the dynamics of such systems Yang2018 ; Wang2020 . The Moiré-like patterns presented in this work are based on a somewhat different mechanism. They mainly arise from the quasi periodicity resulting from the non-linear potential.
This paper is organized as follows. In Section II, the non-Hermitian SSH model under a position-dependent potential is introduced. We explore the symmetries of the Hamiltonian with a linear potential and investigate the possible structures of energy levels based on these symmetries. In Section III and Section IV, we employ the non-Hermitian Floquet method to calculate the phase diagram. In Section V, the dynamical behavior for the case with a linear potential is studied. We also propose a model with non-linear potentials and demonstrate the Moiré-like pattern in the dynamics through numerical simulations. A summary is provided in the Section VI.
Figure 1: Schematic illustrations of (a) the non-Hermitian SSH Hamiltonian with a tilted field in Eq. (1), which consists of adjacent real hopping strength $J$ and imaginary hopping strength $i\beta$, and (b-d) its consponding energy level structures for three typical cases. When the on-site tilted potential has a uniform slope $\omega$, the Hamiltonian obeys the three symmetries defined by Eqs. (2), (5), and (10), respectively. These constrain the energy spectrum partition the energy spectrum into three distinct categories: (b) two real ladders with an energy shift, forming a symmetric spectrum with respect to zero energy; (c) a single coalescing ladder with level spacing $2\omega$, (d) two complex conjugate pair ladders. The black solid lines represent the real part of the energy levels, while the pink and blue blocks represent the positive imaginary part and negative imaginary part of the energy levels, respectively.
II Hamiltonian and EP Wannier Stark ladder
We consider a non-Hermitian variant of SSH chain Su1980 ; Takayama1980 ; Roth1987 ; Jackiw1976 ; Heeger1988 ; Ryu2002 ; Hasan2010 ; Qi2011 imposed by a position dependent field with the Hamiltonian
$H$ $=$ $\sum\limits_{j=-\infty}^{+\infty}[J(|2j\rangle\langle 2j+1|+\,\mathrm{h.c.})+i\beta(|2j\rangle\langle 2j-1|$ (1)
$+\,\mathrm{h.c.})]+\sum\limits_{l=-\infty}^{+\infty}V_{l}|l\rangle\langle l|]\,,$
where $|l\rangle$ denotes a site state describing the Wannier state localized on the $l$th period of the potential. Here, $J$ and $i\beta$ are the tunneling strength, and $V_{l}$ a on-site potential. In this paper, we take $J$ and $\beta>0$ for the sake of simplicity. The system is schematically illustrated in Fig. 1(a). In previous work, it has been shown that the energy levels obeys a simple structures in its real and imaginary parts when taking the linear potential, $V_{l}=\omega(l+\frac{1}{2})$ Zhang2024 . In this work, we consider the case involving a non-linear potential $V_{l}$, which will be demonstrated to give rise to a dynamic Moiré-like pattern.
To proceed, we start with the case with $V_{l}=\omega(l+\frac{1}{2})$. The structure of energy levels of $H$ can be obtained from its special symmetries without the necessity for detailed calculations. We note that the Hamiltonian has $\mathcal{PT}$ symmetry, satisfying
$\mathcal{PT}H\left(\mathcal{PT}\right)^{-1}=H,$ (2)
where the linear operator $\mathcal{P}$ and antilinear operator $\mathcal{T}$ are defined as
$\left\{\begin{array}[]{c}\mathcal{P}|2{l}\mathcal{\rangle}=(-1)^{l}|2l\mathcal{\mathcal{\rangle}}\\ \mathcal{P}|2l+1\mathcal{\mathcal{\rangle}}=(-1)^{l}|2l+1\mathcal{\mathcal{\rangle}}\end{array}\right.,\,$ (3)
and $\mathcal{T}i\mathcal{T}^{-1}=-i$. According to non-Hermitian quantum mechanics Scholtz1992 , a $\mathcal{PT}$-symmetric Hamiltonian, as a pseudo-Hermitian system, is characterized by energy levels that appear either real or in complex conjugate pairs. The eigenstates corresponding to complex conjugate energy levels are interconnected by the $\mathcal{PT}$ operation. Introducing the translation operator $T_{2}$ defined as
$T_{2}|l\rangle=|l+2\rangle,$ (4)
the special structure of the linear potential ensures that $H$ obeys
$T_{2}HT_{2}^{-1}=H-2\omega,$ (5)
which is referred to as a ramped translational symmetry. Suppose we have a solution $|\psi_{0}^{+}\rangle$ of the Schrodinger equation corresponding to energy $E_{0}^{+}$
$H|\psi_{0}^{+}\rangle=E_{0}^{+}|\psi_{0}^{+}\rangle,$ (6)
a set of eigenstates can be generated as
$|\psi_{n}^{+}\rangle=\left(T_{2}\right)^{n}|\psi_{0}^{+}\rangle$ (7)
with eigen energy $E_{n}^{+}=E_{0}^{+}+2n\mathcal{\omega}$$(n=0,\pm 1,\pm 2,...)$, which covers half of the complete set eigenstates. In paralell, one can construct another set of eigenstates
$|\psi_{n}^{-}\rangle=\left(T_{2}\right)^{n}|\psi_{0}^{-}\rangle$ (8)
with eigen energy $E_{n}^{-}=E_{0}^{-}+2n\mathcal{\omega}$$(n=0,\pm 1,\pm 2,...)$, by starting an eigenstate $|\psi_{0}^{-}\rangle$. Without loss of generality, we can assume that the real part of the energy eigenvalues, $\left|\mathrm{Re}(E_{0}^{\pm})\right|$, is less than or equal to $\mathcal{\omega}$.
We find that the structure of the entire energy spectrum is solely determined by the base energy level $E_{0}^{\pm}$. As a pseudo-Hermitian system, it is straightforward to conclude that its eigenvalues $E_{0}^{+}$ and $E_{0}^{-}$ are either real or form complex conjugate pair. Additionally, there exists another symmetry within the system that provides a constraint for the relationship between $E_{0}^{+}$ and $E_{0}^{-}$. In fact, introducing a modified reflection transformation with respect of the position $m-1/2$,
$R_{m}|l\rangle=(-1)^{l}|2m-1-l\rangle,$ (9)
we have
$R_{m}HR_{m}^{-1}=-H+2m\omega.$ (10)
Applying the above relation on the state $|\psi_{n}^{\pm}\rangle$, we have
$HR_{m}|\psi_{n}^{\pm}\rangle=\left(2m\omega-E_{n}^{\pm}\right)R_{m}|\psi_{n}^{\pm}\rangle.$ (11)
Specifically, taking $m=0$, we have
$H\left(g_{0}|\psi_{0}^{\pm}\rangle\right)=-E_{0}^{\pm}\left(g_{0}|\psi_{0}^{\pm}\rangle\right),$ (12)
which results in nothing but
$E_{0}^{+}=-E_{0}^{-}.$ (13)
Based on the above analysis, we conclude that: (i) When $E_{0}^{+}$ is complex, $E_{0}^{\pm}$ must be imaginary, and the two sets of eigenstates can be generated as
$|\psi_{n}^{+}\rangle=\left(T_{2}\right)^{n}|\psi_{0}^{+}\rangle=\mathcal{PT}|\psi_{n}^{-}\rangle.$ (14)
(ii) When $E_{0}^{+}=E_{0}^{-}=0$, the two sets of WS ladder states coalesce into a single set, that is, $|\psi_{n}^{+}\rangle=|\psi_{n}^{-}\rangle$. (iii) When $E_{0}^{+}$ is real, the two sets of energy levels are real, with an energy shift of $2\left|E_{0}^{+}\right|$, forming a symmetric spectrum with respect to zero energy.
The structure of energy levels for each case is schematically illustrated in Fig. 1(b-d). In the following sections, we will confirm the aforementioned predictions and provide a detailed solution by introducing the Floquet method. We will demonstrate that the EP ladder, which serves as the boundary between complex and real regions, emerges periodically with minor variations in the slope of the linear field. These allow for the appearance of dynamic Moiré-like patterns in the system with a weak non-linear field.
III Non-Hermitian Floquet method
In this section, we investigate the energy level structure within the framework of the non-Hermitian Floquet method, which represents an extension of the conventional Floquet method Maksimov2015 to the presnt non-Hermitian Hamiltonian.
Setting
$|\psi\rangle=\sum\limits_{l=-\infty}^{+\infty}(\phi_{l}^{A}|2l\rangle+\phi_{l}^{B}|2l+1\rangle)$ (15)
the Schrodinger equation $H|\psi\rangle=E|\psi\rangle$ can be written as the form
$2\omega(l+\frac{1}{4})\phi_{l}^{A}-J\phi_{l}^{B}-i\beta\phi_{l-1}^{B}$ $=$ $E\phi_{l}^{A},$ (16)
$2\omega(l+\frac{3}{4})\phi_{l}^{B}-J\phi_{l}^{A}-i\beta\phi_{l+1}^{A}$ $=$ $E\phi_{l}^{B}.$
To solve the coupled equations, we introduce the Bloch wave representation by the Fourier transformation
$\phi_{k}^{A,B}=\frac{1}{\sqrt{2\pi}}\sum_{l=-\infty}^{+\infty}\phi_{l}^{A,B}e^{-ikl},$ (17)
with the wave vector $k\in[-\pi,\pi)$, which represents the Eqs. (16) as two ordinary differential equations
$i2\omega\frac{\mathrm{d}\mathbf{Y}_{k}}{\mathrm{d}k}=G_{k}\mathbf{Y}_{k}.$ (18)
Here the vector $\mathbf{Y}_{k}$ is defined as $\mathbf{Y}_{k}=\left(\phi_{k}^{A},\phi_{k}^{B}\right)^{\intercal}$ and $G_{k}$ is a $2\times 2$ non-Hermitian matrix
$G_{k}=\left(\begin{array}[]{cc}E-\frac{\omega}{2}&-J-i\beta e^{-ik}\\ -J-i\beta e^{ik}&E-\frac{3}{2}\omega\end{array}\right).$ (19)
Owing to the requirement for the wave function to be single-valued, we focus solely on periodic solutions that satisfy $\mathbf{Y}_{2\pi+k}=\mathbf{Y}_{k}$. This leads to the quantization of the energy $E$ in the Eq. (19).
In fact, Eq. (18) can be regarded as a Schroinger equation for a non-Hermitian time-dependent Hamiltonian $G_{k}$, assuming $k$ as time $t$. The periodicity of solutions necessitates that the Floquet operator
$U=\mathcal{T}\exp\left(\frac{1}{2i\omega}\int_{0}^{2\pi}G_{k}\mathrm{d}k\right),$ (20)
has an eigenvalue equal to $1$, where $\mathcal{T}$ is the time-order operator. To proceed, we focus on the $2\times 2$ matrix
$U_{0}=\exp(\frac{iE\pi}{\omega})U=\mathcal{T}\exp\left(\frac{1}{2i\omega}\int_{0}^{2\pi}G_{k}^{0}\mathrm{d}k\right),$ (21)
where the matrix $G_{k}^{0}$ denotes the matrix $G_{k}$ at $E=0$. Then the eigen energy $E_{1,2}$ can be obtained from the two eigen values $\lambda_{1}$ and $\lambda_{2}$ of the matrix $U_{0}$, by the relation
$\lambda_{1,2}=\exp(i\pi E_{1,2}/\omega).$ (22)
For a given pair of ($\lambda_{1}$, $\lambda_{2}$), we always have
$\left\{\begin{array}[]{c}E_{1}=\varepsilon_{1}+2m\mathcal{\omega}\\ E_{2}=\varepsilon_{2}+2n\mathcal{\omega}\end{array}\right.,$ (23)
$(m,n=0,\pm 1,\pm 2,...)$ with
$\lambda_{1,2}=\exp(i\pi\varepsilon_{1,2}/\omega),$ (24)
[…]
In general, analytical expressions for $\lambda_{1,2}$ and $\varepsilon_{1,2}$ cannot be obtained because they involve the time evolution of a time-dependent system. However, it provides a method for solving the coupled differential equations, which serves as the foundation for numerical simulations. Notably, the Floquet method can be employed to investigate the structure of the energy levels in a non-Hermitian Stark ladder system, which is more complicated than that in a Hermitian system. In the subsequent sections, we will delve into the application of this method, exploring its practical implications and revealing the spectacular features of a non-Hermitian SW ladder system.
Figure 2: (a) The phase diagram of the Hamiltonian, as depicted in Eq. (1), illustrating the different phases of the system. The green and blue areas correspond to the parameter ranges where the real non-degenerate energy levels in Fig. 1(b) and the complex conjugate pairs of energy levels in Fig. 1(d) occur, respectively. The red circles represent the distribution of the EP boundary between phases, while the red solid line is the result predicted by the Eq. (39). (b) The same phase diagram as (a) is presented with an additional specific line, and four typical points are labeled. The dotted line represents the parameters with $\beta=0.46J$, varying $J/\omega$ from $9.8$ (point A) to $16.2$ (point D). This dotted line goes through two distinct phases, indicating the transitions from one state to another periodically. The dynamic behaviors for the systems with parameters corresponding to points C and B are presented in Figs. 3(a) and 3(b), respectively. This line provides a way to demonstrate different dynamic behaviors along a chain with the non-linear potential given by Eq. (47), exhibiting Moiré-like patterns as shown in Fig. 4.
IV Phase diagram
IV.1 Phase diagram
Now, we turn to investigate the further relationship between $\varepsilon_{1,2}$ and $\lambda_{1,2}$. The determinant of matrix $U_{0}$ can be obtained as
$\det|U_{0}|=\mathcal{T}\exp[\frac{1}{2i\omega}\int_{0}^{2\pi}\text{tr}(G_{k}^{0})\mathrm{d}k]=1,$ (25)
which leads to $\lambda_{1}\lambda_{2}=\exp[i\pi(\varepsilon_{1}+\varepsilon_{2})/\omega]=1$, or
$\varepsilon_{1}+\varepsilon_{2}=0.$ (26)
There exist three typical cases: (i) In the case with $\lambda_{1}=\lambda_{2}^{\ast}=e^{iv\pi}$ with real $v$, we have $\varepsilon_{1}=-\varepsilon_{2}=v\omega$, and then two sets of real WS ladders. (ii) In the case with $\lambda_{1}=\lambda_{2}^{-1}=e^{v\pi}$, we have $\varepsilon_{1}=-\varepsilon_{2}=-iv\omega$, and then two sets of WS ladders, which are all mutually conjugate. (iii) In the case with $\lambda_{1}=\lambda_{2}=1$, we have $\varepsilon_{1}=\varepsilon_{2}=0$, and then single set of coalescing ladder. We demonstrate this by the following simple examples. Considering $J=0$, $\beta\neq 0$, and $\omega\neq 0$, the system simplifies to a set of independent non-Hermitian dimers. It is evident that the energy levels are composed of two real or complex Stark ladders, or they may coalesce into a single ladder. Now we investigate it within the context of the Floquet method. The corresponding matrix $G^{0}_{k}$ is given by
$G^{0}_{k}=-\left(\begin{array}[]{cc}\frac{\omega}{2}&i\beta e^{-ik}\\ i\beta e^{ik}&\frac{3}{2}\omega\end{array}\right),$ (27)
where the phases $e^{\pm ik}$ can be eliminated without changing the physics. Then the matrix $U_{0}$ is obtained as
$U_{0}=\cos\Lambda+\frac{i\pi\sin\Lambda}{\Lambda}\left(\begin{array}[]{cc}\frac{\omega}{2}&\frac{i\beta}{\omega}\\ \frac{i\beta}{\omega}&-\frac{\omega}{2}\end{array}\right),$ (28)
with $\Lambda=\pi\sqrt{\frac{1}{4}-\left(\frac{\beta}{\omega}\right)^{2}}$. The eigenvalues are given by
$\lambda_{1}=\lambda_{2}^{-1}=\exp(i\Lambda),$ (29)
which indicates the following scenarios: (i) there are two sets of real ladders when $\beta<\omega/2$; (ii) there are two sets of complex ladders when $\beta>\omega/2$; and (iii) there is a single set of coalescing ladders when $\beta=\omega/2$. Furthermore, these examples demonstrate that the distinction between a degenerate ladder and a coalescing ladder stems from whether their corresponding matrix $U_{0}$ is a scalar matrix or a Jordan block.
For a system with a broad range of parameters $(J,\beta,\omega)$, the types of Stark ladders in an infinite chain can be ascertained through numerical simulations employing the Floquet method. In Fig. 2(a) we plot the phase diagram in the $J/\omega-\beta/\omega$ plane.
IV.2 The distribution of EPs under the weak-field approximation
The numerical results show that for fixed $(J,\beta)$ the position of EP strongly depends on the value of $\omega$. This is crucial for the present work when we consider the case where the external potential $V_{l}$ is not linear. In the following we will establish the relation among the parameter $(J,\beta,\omega)$ at EP based on the weak-field approximation.
We rewrite the matrix $G_{k}$ in Eq. (19) in the form $G_{k}=h_{0}+h^{\prime}$, where
$h_{0}=\left(\begin{array}[]{cc}E-\frac{\omega}{2}&0\\ 0&E-\frac{3}{2}\omega\end{array}\right)$ (30)
is in the diagonal form, while $h^{\prime}$ is given by
$h^{\prime}=\left(\begin{array}[]{cc}0&-J-i\beta e^{-ik}\\ -J-i\beta e^{ik}&0\end{array}\right).$ (31)
In the spirit of the interaction picture in quantum mechanics, the Floquet operator can be expressed as
$U$ $=$ $\mathcal{T}\exp\left(\frac{1}{2i\omega}\int_{0}^{2\pi}(h_{0}+h_{\mathrm{I}})\mathrm{d}k\right)$ (32)
$=$ $i\exp(-i\frac{\pi E}{\omega})\sigma_{z}\mathcal{T}\exp\left(\frac{1}{2i\omega}\int_{0}^{2\pi}h_{\mathrm{I}}\mathrm{d}k\right),$
where
$h_{\mathrm{I}}$ $=$ $e^{-ih_{0}k}h^{\prime}e^{ih_{0}k}$ (35)
$=$ $-\left(\begin{array}[]{cc}0&Je^{ik/2}+i\beta e^{-ik/2}\\ Je^{-ik/2}+i\beta e^{ik/2}&0\end{array}\right).$
Then the calculation of $\mathcal{T}\exp\left(\frac{1}{2i\omega}\int_{0}^{2\pi}h_{\mathrm{I}}\mathrm{d}k\right)$ can be implemented by the time evolution of the equivalent Schrodinger equation
$i\frac{\mathrm{d}\mathbf{\Phi}(t)}{\mathrm{d}t}=-\left(\begin{array}[]{cc}0&Je^{i\omega t}+i\beta e^{-i\omega t}\\ Je^{-i\omega t}+i\beta e^{i\omega t}&0\end{array}\right)\mathbf{\Phi}(t),$ (36)
where $t=k/\left(2\omega\right)$ can be regarded as time, while $\omega$ is the frequency of the varying Hamiltonian. It is still not exactly solvable for a given $\omega$. However, an adiabatic solution is available when the system varies in the limit $\omega\rightarrow 0$. For small $\omega$, the adiabatic approximation corresponds to the weak-field approximation.
For an adiabatic process, the evolved state of an initial eigenstate is the simultaneous eigenstate of the time-dependent Hamiltonian. Specifically, we have
$\mathbf{\Phi}(\pi/\omega)=-\exp i(\varphi_{\mathrm{D}}+\varphi_{\mathrm{B}})\sigma_{z}\mathbf{\Phi}(0),$
where $\mathbf{\Phi}(0)$ can be one of the eigenstates of the initial Hamiltonian. The constraint condition given by
$U=-i\exp(-i\frac{\pi E}{\omega})\exp i(\varphi_{\mathrm{D}}+\varphi_{\mathrm{B}})=1,$ (37)
requires that
$\pi E/\omega+\frac{\pi}{2}=\varphi_{\mathrm{D}}+\varphi_{\mathrm{B}}.$ (38)
A straightforward derivation shows that
$\varphi_{\mathrm{D}}=\frac{d}{\omega},\,\varphi_{\mathrm{B}}=-\frac{\pi}{2},$
where $d=\int_{0}^{\pi}\sqrt{J^{2}-\beta^{2}+i2\beta J\cos\theta}\mathrm{d}\theta$, which can be evaluated by the elliptic integral of the second kind. Together with the relation $E=2n\omega$ at EP, we get the EP curve given by
$\omega=\frac{d}{4n\pi+2\pi}.$ (39)
In Fig. 2(a), we plot the phase diagram in comparison with the numerical results. It shows that they are consistent with each other for large values of $\left(J/\omega,\beta/\omega\right)$.
Figure 3: The profiles of the evolved states for the initial state given by Eq. (45) under the systems with parameters corresponding to points C and B in Fig. 2(b). 3D plots of (a) $P_{n}(t)$ and (b) $P_{n}/P$, as defined in Eqs. (43) and (46), are obtained by exact diagonalization for finite system with $N=100$. We observe that the evolved state (a) is a superposition of two Bloch oscillations with a period $T=\pi/\omega$, and (b) represents the amplified Bloch oscillations as expected. The logarithm of the total Dirac probability, $\text{In}P(t)$, indicates that difference between two kinds of dynamic behaviors is evident.
V Dynamics Under Linear and Nonlinear Potentials
Now we turn to the dynamic characteristics of the phase diagram, starting with the equations
$\left\{\begin{array}[]{c}H|\psi_{n}^{+}\rangle=\left(E_{0}^{+}+2n\mathcal{\omega}\right)|\psi_{n}^{+}\rangle\\ H|\psi_{m}^{-}\rangle=\left(-E_{0}^{+}+2m\mathcal{\omega}\right)|\psi_{m}^{-}\rangle\end{array}\right.,$ (40)
[…]
$|\Phi(0)\rangle=\sum_{l,\sigma=\pm}c_{l}^{\sigma}|\psi_{l}^{\sigma}\rangle,$ (41)
the time evolution state is
[…]
where $\left\{|\psi_{l}^{\pm}\rangle\right\}$ is normalized with Dirac inner product by taking $\left|\langle\psi_{l}^{\pm}|\psi_{l}^{\pm}\rangle\right|^{2}=1$. (i) In the case with $E_{0}^{+}=\left|E_{0}^{+}\right|$, it is clear that the evolved state is the superposition of two Bloch oscillations with a period of $\pi/\omega$. We are interested in the Dirac probability of the evolved state, which is defined by
$P(t)=\left|\langle\Phi(t)|\Phi(t)\rangle\right|^{2}.$ (43)
In this case, $P(t)$ is bounded function with a constant average. (ii) In the case with $E_{0}^{+}=i\left|E_{0}^{+}\right|$, we have
$|\Phi(t)\rangle\approx e^{E_{0}^{+}t}\sum_{l}c_{l}^{+}e^{-i2l\mathcal{\omega}t}|\psi_{l}^{+}\rangle,$ (44)
for large $t$. It exhibits amplified Bloch oscillations, characterized by a probability that increases exponentially. To demonstrate this point, we perform numerical simulations for a system with a linear potential. We take a Gaussian wavepacket as the initial state, which is in the form
$\left|\Phi(0)\right\rangle=\frac{1}{\sqrt{\Omega}}\sum\limits_{l=-N/2}^{N/2}\exp(-0.01l^{2})\left|l\right\rangle.$ (45)
where $\Omega$ is the normalization coefficient. The evolved state is $\left|\Phi(t)\right\rangle=e^{-iHt}\left|\Phi(0)\right\rangle$, which can be obtained by numerical simulation for a finite size system. The time-dependent Dirac probability distribution is
$P_{n}(t)=\left|\left\langle n\right|\Phi(t)\rangle\right|^{2}.$ (46)
The numerical results with two typical sets of parameters are plotted in Fig. 3. As predicted, the evolved state exhibits the superposition of two Bloch oscillations with a period of $\pi/\omega$ when the system has a pure real spectrum. In contrast, the evolved state exhibits amplified Bloch oscillations when the system has a full complex spectrum. The evident difference between two kinds of dynamic behaviors lies in the total Dirac probability: the former has a constant average, while the latter grows exponentially as time increases.
Figure 4: The profile of the evolved state, given the initial state in Eq. (48), is computed for the system with $\beta=0.46J$ ($J=1$) and a non-linear potential as described by Eq. (47). The 3D plot of $P_{n}(t)$, as defined in Eq. (46), shows that the evolving state exhibits evident Moiré-like patterns. The paremeters are $T=\pi\tau=13\pi$,$\ \gamma=1.6\times 10^{-3}$, and $N=4000$.
We now know that in a system with fixed $J$ and $\beta$, the EP appears as $\omega$ varies, which is associated with distinct dynamic behaviors for a given initial state. It indicates that the EPs occur at different locations when $\omega$ varies slowly along the chain. Accordingly, the local dynamics should also be position-dependent. Our primary interest here is in the effect of slowly varying position-dependent nonlinear potentials on system dynamics. The potential can map the periodic structure of EPs to the local dynamics behavior, and induce the formation of a Moiré-like pattern. To demonstrate this point, we perform numerical simulations for a system with a non-linear potential of the form
$V_{l}=\frac{1}{\gamma l+\tau}(l+\frac{1}{2}),$ (47)
where the constants $\gamma$ and $\tau$ satisfy $\gamma\ll 1/J\ll\tau$. It allows mapping the parameter line in Fig. 2(b) onto an SSH chain, varying the local $J/\omega$ from $9.8$ to $16.2$ along the chain. We take the initial state being distributed on each site with the equal probability
$\left|\Phi(0)\right\rangle=\frac{1}{\sqrt{N}}\sum\limits_{l=-N/2}^{N/2}\left|l\right\rangle.$ (48)
We plot the probability distributions $P_{n}(t)$ in Fig. 4. It can be observed that the evolving state exhibits apparent Moiré-like patterns. Before ending this paper, it is worth briefly mentioning that the Moiré-like pattern previously discussed can be implemented through an array of evanescently coupled optical waveguides Makris2008 ; Musslimani2008 ; Klaiman2008 ; Longhi2009 ; ElGanainy2007 . In this context, the time evolution of state $\left|\Phi(t)\right\rangle$ mapped into the spatial evolution of the modal amplitudes of light waves, along the array axis.
VI Summary
In summary, we have investigated the dynamical behavior of the non-Hermitian SSH chain, focusing on the effects induced by a non-linear field rather than a linear one. We have shown that, by analytical analysis and numerical simulation based on the non-Hermitian Floquet method, such a system with a linear potential exhibits the following features: (i) The spectrum of the system consists of two sets of complex WS ladders; (ii) they can be two real WS ladders with identical level spacing, a single coalescing WS ladders referred to as an EP ladder, and complex conjugate piar of WS ladders; (iii) As the boundary between the complex and real regions, the EP ladder appears periodically as the slope of the linear field changes slightly. These allow for the appearance of dynamic Moiré-like patterns in the system with a weak non-linear field. Our findings not only shed light on the dynamical properties of non-Hermitian Hamiltonians but also offer a platform for implementing novel optical phenomena.
Acknowledgements.
This work was supported by National Natural Science Foundation of China (under Grant No. 12374461).